OpenStudy (anonymous):

Which equation is represented by the table? x y (x, y) 0 3 (0, 3) 1 5 (1, 5) 2 7 (2, 7) A. y = x + 3 B. y = 2x + 3 C. y = 5x D. y = 3x + 2

OpenStudy (anonymous):

plz help

OpenStudy (texaschic101):

pick two sets of points... (0,3) and (1,5) use the slope formula and find the slope slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (5 - 3) / (1 - 0) slope(m) = 2/1 which equals 2. now use y = mx + b slope(m) = 2 you can choose any of the points....I will choose (0,3) now we sub 3 = 2(0) + b 3 = b your equation is : y = 2x + 3